OpenStudy (anonymous):

Explain, in your own words, the step-by-step process to simplifying the expression below. Include the simplified answer in your explanation. (6x2 − 5) − (3x2 + x + 7)

OpenStudy (welshfella):

the first thing you do is remove the 2 parentheses removing the first one makes no difference to the contents. But removing the second one does. You have to distribute the negative over the contents of the second one. Its like multiplying the contents by -1. so we have (6x2 − 5) − (3x2 + x + 7) = 6x^2 - 5 -3x^2 -x - 7 Note that -1 * x = -x and -1*+7 = -7

OpenStudy (welshfella):

Now you simplify like terms to do this you can bring the like terms together (though this is not essential) 6x^2 - 5 -3x^2 -x - 7 = 6x^2 - 3x^2 - x - 5 - 7 = 3x^2 - x - 12

OpenStudy (welshfella):

what is left is called a trinomial or a quadratic which can sometimes be factored . This one cannot be factored.

OpenStudy (welshfella):

So this is the simplified form.

OpenStudy (anonymous):

So the simplified form is 3x^2-x-12?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

Thank you

OpenStudy (welshfella):

yw